Calculate the force between the two point charges q1 = 2.25 μc and q2 = 1.50 μc if they are 10 cm apart.


1
Expert's answer
2021-05-10T02:42:32-0400

Force of attraction or repulsion is given by coulomb law

F = k|q1| |q2|/r^2

where k = 8.9876 × 10^9 N·m^2/C^2

r= distance between two points charge=10cm=0.10m

q1=2.25μC=2.25*10^(-6)C

q2=1.5μC=1.5*10^(-6)C

F=k|q1| |q2|/r^2

F=8.9876 × 10^9x2.25x10^(-6)x1.5*10^(-6)/(0.10)^2

F=3.033N